OpenStudy (anonymous):

can someone help me how to show that there is no rational number x such that xsquared =2

OpenStudy (zzr0ck3r):

assume by contradiction sqrt(2) is rational then sqrt(2) = q/p where q and p are relative prime integers (they are reduced), and p is not 0. square both sides 2=p^2/q^2 2q^2=p^2 so p^2 is even (it is not hard to show that p is then even) so p is even so 2q^2=(2k)^2 for some k in integer so 2q^2=4k^2 q^2=2k^2 so q is even so p and q are even this is a contradiction because (p,q) = 1 thus sqrt(2) is irrational

OpenStudy (zzr0ck3r):

if you need me to show that when a^2 is even a must be even I will

OpenStudy (blurbendy):

(p,q) = 1 because they are the same even integer?

OpenStudy (zzr0ck3r):

no (p,q) = 1 means the are relatively prime, they share no common factor.

OpenStudy (zzr0ck3r):

in other words p/q is in "lowest terms" so p and q cant both be even, because then they are not in lowest terms

OpenStudy (zzr0ck3r):

and by the fundamental theorem of arithmetic we know that every ratio can be written in lowest terms....(but this is obvious).

OpenStudy (zzr0ck3r):

understand @gachauJM @blurbendy

OpenStudy (zzr0ck3r):

to show a^2 implies a is even assume a^2 is even but a is odd then a = 2k+1 for some integer k so a^2=(2k+1)^2 = 4k^2+4k+1 = 2(2k^2+2k)+1 and since 2k^2+2k is an integer, call it z, we have a^2 =2z+1 which is of course odd, but a^2 is even, and thus again by contradiction a must be even

OpenStudy (anonymous):

@zzrOck3r i got it.thanks